Attendees: Bruce Glagola, Nena Moonier, Clay White, Bruno Fieramosca, and Patti Pedergnana
- G. Srajer emailed Bruce about user notification of the LNDS upgrade during shutdown, 4/28-5/7. There is a draft email written to notify our users. The LN2 fill stations will continue to operate during this time. LN2 fill station demand information is needed. A survey will be sent with the email notification. Responses will be sent to email@example.com. Patti will compile the information as it comes in.
- On April 22nd there will be a machine shop walkthrough. Clay will join the team on this walk through.
- Today is Tech Tuesday; Clay will attend 433, Bruno 431, and Patti 438.
- The Remote Shutter Interface at 16-BM-D was replaced by the SI group this morning.
- The 434 LOM and beamlines will have their routine ESH walkthroughs today from 1330-1500.
- Sector 18 had the Rossman group BSL-2 experiment about ten days ago. A capillary was broke immediately, and the holder was decontaminated. The beamline area and holders were decontaminated again at the end of the experiment.
- The clean sweep of the gas yards is underway. It will take several weeks to rid of the legacy cylinders.